OpenStudy (anonymous):

What are the period and phase shift for f(x) = tan(2x + pi)?

OpenStudy (aum):

If f(x) = tan(Ax + B), then: Period = pi / A Phase shift = -B/A

OpenStudy (anonymous):

-pi/2?

OpenStudy (aum):

That is the phase shift.

OpenStudy (anonymous):

and pi/2?

OpenStudy (aum):

Yes, Period = pi/2

OpenStudy (anonymous):

and can yu help me with What are the sine, cosine, and tangent of 5 pi over 4 radians?

OpenStudy (anonymous):

you are the best btw

OpenStudy (anonymous):

sin Θ = square root 2 over 2; cos Θ = negative square root 2 over 2; tan Θ = −1 sin Θ = negative square root 2 over 2; cos Θ = negative square root 2 over 2; tan Θ = 1 sin Θ = square root 2 over 2; cos Θ = negative square root 2 over 2; tan Θ = 1 sin Θ = negative square root 2 over 2; cos Θ = square root 2 over 2; tan Θ = −1 answer choices

OpenStudy (aum):

In which quadrant is the angle 5pi/4 radians?

OpenStudy (anonymous):

the 3rd

OpenStudy (aum):

Correct. What are the signs of sine, cosine and tangent in the 3rd quadrant?

OpenStudy (anonymous):

-2√2,-2√2, 1? idrk tangent im pretty sure thats it

OpenStudy (aum):

Yes.

OpenStudy (anonymous):

yay thnk you! i have like 7 more can you help real quick please :*

OpenStudy (aum):

tan = sin / cos. So if you know sine and cosine as you listed above you can easily find tan by dividing: sin / cos. Just for variety let someone else help you while I help someone else. that way you will get a different perspective on doing the problems and I will have some variety answering too.

OpenStudy (anonymous):

ok :( no one else is helping me though